# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 5.4. Quasi-uniformization results
In this subsection, we prove Corollary 1.9 and Corollary 1.10 and Theorem 1.11.
@@
Let $X$ be a compact Kähler klt space of dimension $n$ with $K_{X}$ numerically trivial. Since $\mathcal{T}_{X}$ is generically nef with respect to Kähler form $\omega_{0},\cdots,\omega_{n-2}$ by [53, Proposition 6.5], $\mathcal{T}_{X}$ is semi-stable with respect to any Kähler classes $\omega_{0},\cdots,\omega_{n-2}$ since $c_{1}(K_{X})=0$. Then (1.5) follows from the Bogomolov inequality (3.1).
Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover. Note that $\mathcal{T}_{Y}=\gamma^{[*]}\mathcal{T}_{X}$ by the reflexivity. Since $\widehat{c}_{1}(\mathcal{T}_{X})=0$, it follows from Theorem 1.12 that the equality in (1.5) holds if and only if $\mathcal{T}_{Y}$ is a locally free and flat sheaf, which is equivalent to that $Y$ is smooth by the Lipman-Zariski conjecture for klt spaces (cf. [63, Theorem 1.13]) and so is a torus thanks to Yau’s theorem [109]. The proof is complete. ∎
@@
The proof is similar to [55, Proposition 2.20]. Let $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{{X_{\mathrm{reg}}}})$ be a semi-stable reflexive Higgs sheaves. Then [110, Corollary 1.3] implies that $\mathcal{G}_{X_{\mathrm{reg}}}:=\operatorname{End}(\mathcal{E}_{X_{\mathrm{reg}}})$ with the induced Higgs field $\theta_{\mathcal{G}_{X_{\mathrm{reg}}}}$ is semi-stable. The inequality (1.6) has shown in Theorem 3.2. Recall (3.13) that $\mathcal{E}_{X}$ satisfies the equality in (1.3) if and only if $\mathcal{G}_{X}$ satisfies $\widehat{c}_{1}(\mathcal{G}_{X})=0$ and (1.3). Building on Theorem 1.12, we complete the proof. ∎
Let us outline the proof of Theorem 1.11 by building on the results in Sections 3 and 4, following the spirits of [53] and [58, Section 4]. We refer to [53, Section 3.2 and 4.2] for the precise definitions of stability and of the intersection number taken with respect to the non‑pluripolar product.
@@
Let $X$ be a projective klt variety of dimension $n$ with $K_{X}$ big. There is a natural Higgs sheaf on $X$ consists of the reflexive sheaf $\mathcal{E}_{X}:=\Omega_{X}^{[1]}\oplus\mathcal{O}_{X}$ and the Higgs field $\theta_{X}$ defined by
@@
Since $K_{X}$ is big, $X$ admits a canonical model $X_{\text{can}}$ by [4, Theorem 1.2] and denote $h:X\dashrightarrow X_{\text{can}}$ be the birational map induced by the $K_{X}$-MMP. Let $p_{1}:X_{1}\rightarrow X$ and $q_{1}:X_{1}\rightarrow X_{\text{can}}$ be resolutions of $f$. Then there exists an $q_{1}$-effective divisor $E$ such that
@@
and $p_{1}(\mathrm{supp}(E))$ coincides with the $h$-exceptional divisor. By [21, Lemma A.5] (see also [53, Proposition 3.15]), we have
@@
Let us recall the definition of $\widehat{\Delta}(\mathcal{E}_{X})\cdot\langle K_{X}^{n-2}\rangle$ introduced in [53, Section 4.2.1]. We take an orbifold modification $f^{\prime}:Y^{\prime}\rightarrow X$ such that $Y^{\prime}$ admits a standard orbifold structure $Y_{\mathrm{orb}}^{\prime}:=\{U^{\prime}_{\mu},G_{\mu}^{\prime},\pi_{\mu}^{\prime}\}$. $\mathcal{E}_{Y^{\prime}}:=f^{\prime*}\mathcal{E}_{X}/\mathrm{tor}$ induces a torsion-free orbi-sheaf $\mathcal{E}_{\mathrm{orb}}^{\prime}:=\{\widehat{\pi}_{\mu}^{*}\mathcal{F}_{Y^{\prime}}/\mathrm{tor}\}$. According to Lemma 2.2, there exists a projective bimeromorphic morphism $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow Y_{\mathrm{orb}}^{\prime}$ from an effective complex orbifold $Y_{\mathrm{orb}}=\{U_{\mu},G_{\mu},\pi_{\mu}\}$ to $Y_{\mathrm{orb}}^{\prime}$ such that the torsion-free pullback of $\mathcal{E}_{\mathrm{orb}}^{\prime}$ is an orbi-bundle $E_{\mathrm{orb}}^{\prime}$. Set $f:=f^{\prime}\circ g$ and it follows from [53, Proposition 4.18]
@@
Note that on every orbifold chart $U_{\mu}$, $(f\circ\pi_{\mu})^{*}\mathcal{E}_{X}$ admits a functorial Higgs field $(f\circ\pi_{\mu})^{*}\theta_{X}$ by [63, Theorem 1.10] (cf. [37, Construction 5.5]), which induces a Higgs field $\theta_{\mathrm{orb}}^{\prime}$ of $E_{\mathrm{orb}}^{\prime}$ because $E_{\mathrm{orb}}^{\prime}=\{(f\circ\pi_{\mu})^{*}\mathcal{E}_{X}/\mathrm{tor}\}$.
Choose resolutions $p_{2}:X_{2}\to Y$ and $q_{2}:X_{2}\to X_{1}$ of the composition $p^{-1}\circ f:Y\dashrightarrow X_{1}$. Let $V_{\mu}$ be a functorial resolution of singularities of the normalization of $U_{\mu}\times_{Y}X_{2}$; then the group $G_{\mu}$ lifts to a subgroup $\widehat{G}_{\mu}\subset\operatorname{Aut}(V_{\mu})$. Denote by $\widehat{\pi}_{\mu}:V_{\mu}\to V_{\mu}/\widehat{G}_{\mu}$ the quotient map. The triple $\{V_{\mu},\widehat{G}_{\mu},\widehat{\pi}_{\mu}\}$ forms a complex effective orbifold $Z_{\mathrm{orb}}$; we write $Z$ for its underlying complex space, $g_{1}:Z\rightarrow Y$ and $g_{2}:Z\rightarrow X$ for the induced bimeromorphisms. All these morphisms are summarised in the following diagram:
@@
Denote $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ be the pull-back of the Higgs orbi-bundle $(E_{\mathrm{orb}}^{\prime},\theta_{\mathrm{orb}}^{\prime})$ by $g_{1}$. Note that (5.13) implies
@@
Set $p:=f\circ g_{1}:Z\rightarrow X$ and $q:=q_{1}\circ q_{2}\circ g_{2}:Z\rightarrow X_{\text{can}}$. Using the same argument of the proof of [53, Lemma 4.14 (1)], we have that $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ is $\langle g_{1}^{*}f^{*}K_{X}^{n-1}\rangle$-stable and so is $q^{*}K_{\text{can}}^{n-1}$-stable. Furthermore, the Poincare duality holds for orbifolds, we have that $\langle g_{1}^{*}f^{*}K_{X}^{k}\rangle=\pi_{*}\langle\pi^{*}g_{1}^{*}f^{*}K_{X}^{k}\rangle=g_{1}^{*}\langle f^{*}K_{X}^{k}\rangle$ in $H^{2n-4}(Z,\mathbb{R})$ according to the definition of the non-pluripolar product [53, Section 3.1.3], then (5.15) and (5.14) yield
@@
(5.16) $\left(2\widehat{c}_{2}(X)-\frac{n}{n+1}\widehat{c}_{1}^{2}(X)\right)\cdot\langle K_{X}^{n-2}\rangle=\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}})\right)\cdot q^{*}K_{X}^{n-2}.$
@@
Recall [40, Theorem A] that the reflexive cotangent sheaf $\Omega_{X_{\text{can}}}^{[1]}$ is $K_{X_{\text{can}}}^{n-1}$-semistable, which implies that the natural Higgs sheaf $(\mathcal{E}_{X_{\text{can}}},\theta_{X_{\text{can}}})$ on $X_{\text{can}}$ is $K_{X_{\text{can}}}^{n-1}$-stable (cf. [55, Proposition 2.8]). According to the above construction, $q:Z\rightarrow X_{\text{can}}$ satisfies all properties listed in Proposition 3.4 by substituting $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$= $(\mathcal{E}_{X_{\text{can}}},\theta_{X_{\text{can}}})|_{(X_{\text{can}})_{\mathrm{reg}}}$ and $\omega_{0}=\cdots=\omega_{n-2}=\omega$, where $\omega$ is a Kähler form in the class $K_{X}$. Under the assumption that the equality in (1.8) holds, applying Theorem 3.5 and a similar computation of Section 3.1.3 give
@@
(5.17) $0=4\pi^{2}\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}})\right)\cdot q^{*}K_{X}^{n-2}\geq\int_{X_{\text{can}}\setminus\Sigma}\|F_{H_{\infty},\theta_{\text{can}}}^{\perp}\|^{2}_{H_{\infty},\omega}\cdot\frac{\omega^{n}}{n!}$
@@
and so $F_{H_{\infty},\theta_{\text{can}}}^{\perp}\equiv 0$ on $X\setminus Z$. In particular, $\mathcal{G}_{X_{\text{can}}}:=\operatorname{End}(\mathcal{E}_{X_{\text{can}}})$ with the induced Higgs field $\theta_{\mathcal{G}}$ admits a pluri-harmonic metric on $X_{\text{can}}\setminus\Sigma$ and so admits a pluri-harmonic metric on the regular locus of $X_{\text{can}}$ by Proposition 4.8. Then it follows from Proposition 4.14 that for a maximally quasi-étale cover $\gamma:Y\rightarrow X_{\text{can}}$, $\gamma^{[*]}\mathcal{G}_{X_{\text{can}}}$ is locally free and its all Chern classes vanish. By the reflexivity, $\gamma^{[*]}\mathcal{G}_{X_{\text{can}}}=\operatorname{End}(\Omega_{Y}^{[1]}\oplus\mathcal{O}_{Y})$. As $\mathcal{T}_{Y}$ is a direct summand of $\operatorname{End}(\Omega_{Y}^{[1]}\oplus\mathcal{O}_{Y})$, it follows that $\mathcal{T}_{Y}$ is locally free. Therefore $Y$ is smooth by the solution of Lipman-Zariski conjecture for klt spaces [33, 63] and $Y$ satisfies the Miyaoka-Yau equality with respect to $K_{Y}^{n-2}$. Since $K_{Y}=\gamma^{*}K_{X}$ is ample, the proof is complete by the classical result of Yau [108]. ∎
@@
Although [58, Theorem 1.2] assumes that $X$ is smooth in codimension $2$, this condition is in fact unnecessary. The general case follows by combining the argument above with the methods of Section 4 in the same reference.

### v2

@@ 5.4. Quasi-uniformization results
In this subsection, we prove Corollaries 1.9 and 1.10 and Theorem 1.11.
@@
Let $X$ be a compact Kähler klt space of dimension $n$ with $K_{X}$ numerically trivial. Since $\mathcal{T}_{X}$ is generically nef with respect to $(\omega_{0},\cdots,\omega_{n-2})$ by [54, Proposition 6.5] and $c_{1}(K_{X})=0$, $\mathcal{T}_{X}$ is semistable with respect to $(\omega_{0},\cdots,\omega_{n-2})$. Then (1.5) follows from the Bogomolov-Gieseker inequality (3.1).
Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover. We have $\mathcal{T}_{Y}=\gamma^{[*]}\mathcal{T}_{X}$ by reflexivity. Since $\widehat{c}_{1}(\mathcal{T}_{X})=0$, Theorem 1.12 implies that the equality in (1.5) holds if and only if $\mathcal{T}_{Y}$ is locally free and flat. The Lipman-Zariski conjecture for klt spaces (cf. [64, Theorem 1.13]) then yields that $Y$ is smooth. By Yau’s theorem [112], $Y$ is a complex torus and the proof is complete. ∎
@@
The proof is similar to [56, Proposition 2.20]. Let $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{{X_{\mathrm{reg}}}})$ be a $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf. Then [113, Corollary 1.3] implies that $\mathcal{G}_{X_{\mathrm{reg}}}:=\operatorname{End}(\mathcal{E}_{X_{\mathrm{reg}}})$, endowed with the induced Higgs field $\theta_{\mathcal{G}_{X_{\mathrm{reg}}}}$, is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. The inequality (1.6) was established in Theorem 3.2. Recall from (3.13) that $\mathcal{E}_{X}$ satisfies the equality in (1.3) if and only if $\mathcal{G}_{X}$ satisfies $\widehat{c}_{1}(\mathcal{G}_{X})=0$ together with (1.3). The conclusion now follows from Theorem 1.12. ∎
The proof of Theorem 1.11 follows the spirit of [54] and [59, Section 4], building on the results in Sections 3 and 4. For the precise definitions of stability and of the intersection number taken with respect to the non‑pluripolar product, we refer to [54, Section 3.2 and 4.2].
@@
Let $X$ be a projective klt variety of dimension $n$ with $K_{X}$ big. Consider the natural Higgs sheaf on $X$ consisting of the reflexive sheaf $\mathcal{E}_{X}:=\Omega_{X}^{[1]}\oplus\mathcal{O}_{X}$ together with the Higgs field $\theta_{X}$ defined by
@@
Since $K_{X}$ is big, $X$ admits a canonical model $X_{\text{can}}$ (see [4, Theorem 1.2]). Denote by $h:X\dashrightarrow X_{\text{can}}$ the birational map induced by the $K_{X}$-MMP. Let $p_{1}:X_{1}\rightarrow X$ and $q_{1}:X_{1}\rightarrow X_{\text{can}}$ be resolutions of $f$. Then there exists a $q_{1}$-effective divisor $E$ such that
@@
and $p_{1}(\mathrm{supp}(E))$ coincides with the $h$-exceptional divisor. By [21, Lemma A.5] (see also [54, Proposition 3.15]), we have
@@
We now recall the definition of $\widehat{\Delta}(\mathcal{E}_{X})\cdot\langle K_{X}^{n-2}\rangle$ introduced in [54, Section 4.2.1]. Take an orbifold modification $f^{\prime}:Y^{\prime}\rightarrow X$ such that $Y^{\prime}$ admits a standard orbifold structure $Y_{\mathrm{orb}}^{\prime}:=\{(U^{\prime}_{\alpha},G_{\alpha}^{\prime},\mu_{\alpha}^{\prime})\}$. Set $\mathcal{E}_{Y^{\prime}}:=f^{\prime*}\mathcal{E}_{X}/\mathrm{tor}$. This induces a torsion-free orbi-sheaf $\mathcal{E}_{\mathrm{orb}}^{\prime}:=\{\widehat{\mu}_{\alpha}^{*}\mathcal{F}_{Y^{\prime}}/\mathrm{torsion}\}$. According to Lemma 2.2, there exists a projective bimeromorphic morphism $g_{\mathrm{orb}}:Y_{\mathrm{orb}}\rightarrow Y_{\mathrm{orb}}^{\prime}$ from an effective complex orbifold $Y_{\mathrm{orb}}=\{(U_{\alpha},G_{\alpha},\mu_{\alpha})\}$ to $Y_{\mathrm{orb}}^{\prime}$ such that the torsion-free pullback of $\mathcal{E}_{\mathrm{orb}}^{\prime}$ is an orbi-bundle $E_{\mathrm{orb}}^{\prime}$. Set $f:=f^{\prime}\circ g$ and it follows from [54, Proposition 4.18] that
@@
Observe that on every orbifold chart $U_{\alpha}$, $(f\circ\mu_{\alpha})^{*}\mathcal{E}_{X}$ carries a functorial Higgs field $(f\circ\mu_{\alpha})^{*}\theta_{X}$ by [64, Theorem 1.10] (cf. [37, Construction 5.5]), which induces a Higgs field $\theta_{\mathrm{orb}}^{\prime}$ on $E_{\mathrm{orb}}^{\prime}$.
Choose resolutions $p_{2}:X_{2}\to Y$ and $q_{2}:X_{2}\to X_{1}$ of the composition $p^{-1}\circ f:Y\dashrightarrow X_{1}$. Let $V_{\alpha}$ be a functorial resolution of singularities of the normalization of $U_{\mu}\times_{Y}X_{2}$; then the group $G_{\alpha}$ lifts to a subgroup $\widehat{G}_{\alpha}\subset\operatorname{Aut}(V_{\alpha})$. Denote by $\widehat{\mu}_{\alpha}:V_{\alpha}\to V_{\alpha}/\widehat{G}_{\alpha}$ the quotient map. The triple $\{(V_{\alpha},\widehat{G}_{\alpha},\widehat{\mu}_{\alpha})\}$ forms a complex effective orbifold $Z_{\mathrm{orb}}$; we write $Z$ for its underlying complex space, $g_{1}:Z\rightarrow Y$ and $g_{2}:Z\rightarrow X$ for the induced bimeromorphisms. All these morphisms are summarized in the following diagram:
@@
Let $(E_{\mathrm{orb}},\theta_{\mathrm{orb}})$ be the pullback of the Higgs orbi-bundle $(E_{\mathrm{orb}}^{\prime},\theta_{\mathrm{orb}}^{\prime})$ by $g_{1}$. Note that (5.16) implies
@@
Set $p:=f\circ g_{1}:Z\rightarrow X$ and $q:=q_{1}\circ q_{2}\circ g_{2}:Z\rightarrow X_{\text{can}}$. Since Poincaré duality holds for orbifolds, we have $\langle g_{1}^{*}f^{*}K_{X}^{k}\rangle=\pi_{*}\langle\pi^{*}g_{1}^{*}f^{*}K_{X}^{k}\rangle=g_{1}^{*}\langle f^{*}K_{X}^{k}\rangle$ in $H^{2n-4}(Z,\mathbb{R})$ according to the definition of the non-pluripolar product [54, Section 3.1.3]. Then (5.18) and (5.17) yield
@@
(5.19) $\left(2\widehat{c}_{2}(X)-\frac{n}{n+1}\widehat{c}_{1}^{2}(X)\right)\cdot\langle K_{X}^{n-2}\rangle=\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}})\right)\cdot\langle q^{*}K_{X}^{n-2}\rangle.$
@@
Recall [40, Theorem A] that the reflexive cotangent sheaf $\Omega_{X_{\text{can}}}^{[1]}$ is $K_{X_{\text{can}}}^{n-1}$-semistable, which implies that the natural Higgs sheaf $(\mathcal{E}_{X_{\text{can}}},\theta_{X_{\text{can}}})$ on $X_{\text{can}}$ is $K_{X_{\text{can}}}^{n-1}$-stable (cf. [56, Proposition 2.8]). According to the above construction, $q:Z\rightarrow X_{\text{can}}$ satisfies all properties listed in Proposition 3.4 upon substituting $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{X_{\mathrm{reg}}})$= $(\mathcal{E}_{X_{\text{can}}},\theta_{X_{\text{can}}})|_{(X_{\text{can}})_{\mathrm{reg}}}$ and $\omega_{0}=\cdots=\omega_{n-2}=\omega$, where $\omega$ is a Kähler form in the class $K_{X}$. Under the assumption that the equality in (1.8) holds, applying Theorem 3.5 and a computation similar to that in Section 3.1.3 give
@@
(5.20) $0=4\pi^{2}\left(2c_{2}^{\mathrm{orb}}(E_{\mathrm{orb}})-\frac{n}{n+1}({c}_{1}^{\mathrm{orb}})^{2}(E_{\mathrm{orb}})\right)\cdot\langle q^{*}K_{X}^{n-2}\rangle\geq\int_{X_{\text{can}}\setminus\Sigma}\|F_{H_{\infty},\theta_{\text{can}}}^{\perp}\|^{2}_{H_{\infty},\omega}\cdot\frac{\omega^{n}}{n!}$
@@
and hence $F_{H_{\infty},\theta_{\text{can}}}^{\perp}\equiv 0$ on $X_{\text{can}}\setminus\Sigma$. In particular, $\mathcal{G}_{X_{\text{can}}}:=\operatorname{End}(\mathcal{E}_{X_{\text{can}}})$ together with the induced Higgs field $\theta_{\mathcal{G}}$ admits a pluri-harmonic metric on $X_{\text{can}}\setminus\Sigma$. This Higgs sheaf therefore admits a pluri-harmonic metric on the regular locus of $X_{\text{can}}$ by Proposition 4.8. Then applying Proposition 4.14, we obtain that for a maximally quasi-étale cover $\gamma:W\rightarrow X_{\text{can}}$, $\gamma^{[*]}\mathcal{G}_{X_{\text{can}}}$ is locally free and all its Chern classes vanish. By reflexivity, $\gamma^{[*]}\mathcal{G}_{X_{\text{can}}}=\operatorname{End}(\Omega_{W}^{[1]}\oplus\mathcal{O}_{W})$. As $\mathcal{T}_{W}$ is a direct summand of $\operatorname{End}(\Omega_{W}^{[1]}\oplus\mathcal{O}_{W})$, it follows that $\mathcal{T}_{W}$ is locally free. The solution of Lipman-Zariski conjecture for klt spaces [33, 64] then implies that $W$ is smooth. Moreover, $W$ satisfies the Miyaoka-Yau equality with respect to $K_{W}^{n-2}$. Since $K_{W}=\gamma^{*}K_{X_{\text{can}}}$ is ample, the proof is complete by the classical result of Yau [111]. ∎
@@
Although [59, Theorem 1.2] assume that $X$ is smooth in codimension $2$, this condition is in fact unnecessary. The general case follows by combining the argument above with the methods of [59, Section 4].